OpenStudy (anonymous):

how would u answer this =100 plus the product of a number and -2 equals 50. What is the number?

OpenStudy (anonymous):

100 + (x * -2) = 50

OpenStudy (anonymous):

where is the number you are trying to solve. Simply solve for x

OpenStudy (anonymous):

100 + (x * -2) = 50 (x * -2) = 50 - 100 That's how I would start